# Impact of Guidance and Interaction Strategies for LLM Use on Learner Performance and Perception

### v2

@@ Abstract.
large language models, tutoring systems, Human-AI collaboration, artificial intelligence in education, collaborative learning with AI.
journalyear: 2023doi: XXXXXXX.XXXXXXXccs: Social and professional topics Computing educationccs: Human-centered computing Empirical studies in interaction designccs: Human-centered computing Interaction techniques
@@ 1. Introduction
To address this issue, previous research on Intelligent Tutoring Systems (ITS) and Intelligent Teaching Assistants (ITA) has pioneered adaptive learning experiences and personalized instructional feedback (Weitekamp et al., 2020; Anderson et al., 1995), increasing the interest of CSCW community to the role that conversational agents can play in supporting the education process (Richards Maldonado et al., 2023; Wu et al., 2021; Breideband et al., 2023; Wardrip et al., 2013; Liu et al., 2020; Sobel et al., 2017; Davis et al., 2018). The rapid emergence of large language models (LLMs) in recent years has created even more promising opportunities for technological support of students’ personalized attention needs (Zhou et al., 2022; Sejnowski, 2023). These models, with their ability to comprehend and generate human-like text, have the potential to reshape the landscape of personalized education. Consequently, researchers started exploring the potential advantages of the development of LLM-based agent tutors as potential surrogates for human teaching assistants (Rahman and Watanobe, 2023; Dergaa et al., 2023; Baidoo-Anu and Owusu Ansah, 2023; Kumar et al., 2023). Indeed, the dialogic nature of LLM-based agents, their ability to adapt to individualized learning patterns, and their ease of integration into diverse educational contexts make them a potential boon for educational frameworks (Farrokhnia et al., 2023). While learning environments vary between subjects and educational levels, the versatility of LLMs allows their integration into these specialized contexts, providing a unified platform for varied educational needs.
@@
However, although there is a considerable volume of existing research on LLM applications, a noticeable gap remains when applied to education. For instance, an extensive body of pedagogical research (Keil and Wilson, 2000; Borko et al., 1997; Shepard, 2001; Van Driel et al., 2001; Chi et al., 2001) suggests that the success of education and tutoring is significantly defined by the communication strategies, employed by an instructor. Thus, the pedagogical and interaction effects are inherently interconnected, and discerning their nuanced impacts is essential for the effective design and implementation of LLM tutors. However, while LLM-based systems can support different interaction strategies (Lee et al., 2022; Kasneci et al., 2023), it remains unclear, which of these strategies are most beneficial for an agent to assume when interacting with students. For instance, how could these strategies influence students’ reception of the assistance, self-assessment, and their subsequent performance? The matter is further complicated by the evidence that the dynamics of social interactions might differ for human-human and human-agent communications (Doyle et al., 2019; Clark et al., 2019), making the direct integration of pedagogical knowledge on learning patterns into agent behavior even more challenging. Correspondingly, while there is a substantial body of literature on the impact of human-human teaching interaction strategies on learning outcomes, our understanding of how these dynamics translate to an LLM-learner context remains limited. Thus, the question is not only about the feasibility of these LLM-based tutors, but also about the ambiguity surrounding the effects of different interaction strategies on students, and to be able to effectively adapt LLM-based agents for individual tutoring, we first need to understand what interaction strategies should be implemented in these systems for optimal learning support.
@@
What influence do various guidance strategies have on enhancing or diminishing learners’ confidence in their ability to use LLMs for problem-solving at each stage of interaction?
@@
In what specific ways do different guidance methods shape learners’ self-confidence in problem-solving across the various stages of interaction?
@@
To gain a deeper understanding of student interactions with LLMs in educational contexts, in this paper, we distinguish student-agent interaction into three key elements: the Guidance Received, which refers to the directions or suggestions students obtain for using LLMs; the Learner Approach, indicating whether learners first consult the LLM or attempt problem-solving independently; and the LLMs’ Response, emphasizing the importance of dialogue quality with LLM. These components are intertwined, with shifts in one potentially influencing the others. We then empirically investigated these influences through two user studies. In both studies, we explored four variations of the first element, using four distinct guidance strategies: list-based suggestions, example-based instruction, metacognitive questioning (making learners think about their use of LLMs before actual use) and a problem-solving-then-refinement approach (making learners solve a problem first and then refine it with LLMs). These four guidance strategies manifested in two distinct student approaches (second element): directly consulting the LLM or relying on self-first problem-solving. Finally, we manipulated the third element – LLMs’ Response – through pre-interaction manipulations (Brown et al., 2020; Kumar et al., 2022)).
@@
While there is previous research on student-AI interaction, LLMs seem to provide different levels of interaction experience and convincingness, even in the presence of errors. Thus, studying the multidimensional picture of student perceptions and their dynamics and analyzing them in a formative experimental setup, combining preliminary quantitative measures as well as extensive qualitative feedback is needed at this initial stage. Our formative study in an undergraduate computer science classroom ($N=145$) tasked students to solve an assignment while receiving help from an LLM-based chatbot. We explored the impact of our four guidance strategies, and the resulting two approaches, on the student’s performance (in the assignment, as well as in the final exam 2 weeks after the assignment), system perception (helpfulness, confidence in LLM’s response, tolerance of mistakes, willingness to interact again), and their self-assessment (change in their self-efficacy, confidence in their answers). We found a reduction in unrelated student queries to LLM with structured guidance such as List of Suggestions, Metacognitive Questioning, and Solve then refine with LLM approaches, while Example based instruction paradoxically increased such queries. Additionally, certain guidance types promoted deeper engagement with the material, evidenced by fewer instances of verbatim question copying and increased rephrasing, highlighting the impact of guidance on the quality and focus of student-tutor interactions in LLM environments. As we discuss further, from a statistical standpoint, this setting is traditional in its inherent limitations of sample size by course enrollment, student dropout, and selective engagement. It allows us to use it only to gather formative data for different conditions. In addition, exposing students to experimental stimuli also allows for gathering instances of qualitative feedback specific to different stimuli.
@@
Informed by findings from formative deployment, we conducted a controlled study on Prolific (Palan and Schitter, 2018) in which crowdworkers (N=356) were asked to solve math problems at the elementary and high school level with the help of an LLM chatbot. This level of math was selected as it ensures a standardized baseline of complexity and familiarity for a broad adult audience, allowing for a more controlled and consistent assessment of the LLM chatbot’s tutoring capabilities. Besides manipulating the different forms of guidance given to students, the online controlled setting gave us affordances to compare a pre-prompted LLM with an unprompted LLM. Through a 2 (LLM Chatbot: Prompted vs. Unprompted) x 5 (Guidance strategy: G1-List of Suggestions vs. G2-Example-based Instruction vs. G3-Metacognitive Questioning vs. G4-Solve then refine with LLM vs. No Instruction), we explored the impact of interaction strategies on the learners’ performance and their confidence (in LLM’s responses, ability to use LLMs, and their Math-solving ability) across the different stages of interaction (before receiving guidance, immediately after receiving guidance, and after task completion). We did not find statistically significant differences in Learner + AI task performance across different Guidance Types and LLM Types (prompted vs unprompted), we observe interesting patterns in confidence in the ability to use LLMs to solve math problems and trust in LLM responses over different stages of interaction. Guidance shortly boosts confidence in the ability to use LLM, but both measures significantly drop after learners’ attempt to solve real tasks with LLM use. This might suggest a partially negative interaction experience on solving tasks, as well as the need for extra learning support for LLM use in the particular context.
@@ 2.1. Understanding Learner-LLM Interactions in Real-World Settings
With LLM-based conversational systems like ChatGPT (cha, [n. d.]), Bard (bar, [n. d.]), and BingChat (bin, [n. d.]) already out in the wild and used daily by hundreds of thousands of learners, educators and researchers are faced with an urgent need to understand how learners interact with these systems in real-world educational settings. They are both thrilled (Rahman and Watanobe, 2023; Dergaa et al., 2023; Baidoo-Anu and Owusu Ansah, 2023) and concerned (Rahman and Watanobe, 2023; Dergaa et al., 2023) about the intergration of LLM-based chatbots into diverse learning environments as we navigate the uncertainties regarding the impact of this rapidly evolving technology on pedagogy.
The growing body of recent CSCW and HCI literature on incorporating LLM-based conversational agents in education have explored various ways to enhance chatbots to assist learners in different ways, such as using them as persuasive agents to target behavior change through self-reflection (Li et al., 2023), making them more goal-aware (Deng et al., 2023), proactive (Liao et al., 2023), and credible (Candello et al., 2023), and customizing them to take on different roles to facilitate different discussion patterns and system thinking (Nguyen, 2023). However, while such work can provide valuable preliminary insights into how we can design chatbots, they often do not investigate learner interactions in real-world settings. Given that learners are already interacting with widely available and powerful LLMs through tools like ChatGPT, it has become increasingly important to understand how learners are interacting with these tools. In this work, we offer empirical insights of classroom and crowdworker interactions with LLMs from two field studies that contribute to this understanding, and offer design recommendations that can inform the design of future LLM-based chatbot implementations.
@@ 2.2. Leveraging Guidance Strategies to Steer Interactions Toward Learning
In tandem with investigating learner-LLM interactions, we also need to understand how to guide students in ways that support their learning instead of impeding it. Guidance plays a crucial role in helping users navigate how to best interact with chatbots, but the efficacy of specific forms of guidance remains underexplored (Wu et al., 2021).
@@
The CSCW community have begun exploring different facets of this important space. For example, Wu et al. (Wu et al., 2021) studied user preferences for two different guidance types (Example-Based and Rule-Based) at four different guidance timings (Service- Onboarding, Task-Intro, After-Failure, and Upon-Request) based on data from 24 participants, and found that users preferred Example-Based guidance, and wanted to see the guidance as part of the Task-Intro. Zhu et al. (Zhu et al., 2022) built a chatbot prototype that provided task-based instructions to novice workers, and offered results from a pilot study with 7 participants that indicated that back-and-forth conversations might help novice workers follow instructions, and perceive those instructions as being more actionable. Our work contributes to such literature on how different guidance strategies impact learner perceptions and interactions with LLM-based chatbots. A distinguishing feature of our work from existing preliminary studies is that we deployed our experiments two distinct learner groups with comparatively larger sample sizes (145 students in an undergraduate CS classroom, and 356 crowdworkers from Prolific). We study the specific effects of four guidance strategies (providing instructions, presenting examples, using metacognitive questioning, and encouraging initial problem-solving with subsequent refinement)
@@
To inform our selection of guidance strategies, in addition to looking at the types of guidance that have been explored in preliminary work within the CSCW community, we also consulted prior pedagogical literature on various forms of instructional guidance (Bonawitz et al., 2011), and chose to study direct instruction (White, 1988), example-based instruction (Wittwer and Renkl, 2010), and metacognitive strategies (Livingston, 2003; Şen, 2009)—such as posing reflective questions (Ellis et al., 2014; Alt and Raichel, 2020), owing to their documented effectiveness in prior research (Gersten et al., 1986; Adams and Engelmann, 1996; Zendler and Klein, 2018; Anderson et al., 2007; Van Gog and Rummel, 2010; Palincsar, 1986; Bourner, 2003).
@@
To understand why guidance is of particular relevance to LLM-based chatbot design, we can turn to recent research on chatbot interactions of novice AI users that show prompt-writing can be deceptively difficult. The response from LLM-based chatbots is primarily steered by text-based prompts (Wu et al., 2022; Zhou et al., 2022). However, recent research has indicated that crafting effective prompts is challenging (Zamfirescu-Pereira et al., 2023), particularly for those without a deep understanding of AI (Abdellatif et al., 2020; Zamfirescu-Pereira et al., 2023; Fiannaca et al., 2023). The conversation-like interface of these chatbots mimics human interaction (Smestad and Volden, 2019), potentially misleading students into thinking that prompt-writing is straightforward as talking to humans (Hofstadter, 1995). This gap between perceived simplicity and the actual complexity of effective prompt-writing can breed overconfidence. Yet, when the chatbot fails to respond as expected, students’ trust in the AI system (Denny et al., 2023a) can wane, leading to frustration (Pérez et al., 2020). In response to these challenges, informal communities like the r/aipromptprogramming and r/ChatGPT subreddits have emerged to share best practices for prompt-writing. Simultaneously, researchers are starting to establish formal guidelines (Denny et al., 2023b; Liu and Chilton, 2022; Ekin, 2023), but given the nascent nature of this field, much is left to be explored. To enhance student success with LLMs, there is a need to convert these guidelines into a more digestible format, thereby improving LLM literacy (Zamfirescu-Pereira et al., 2023; Yang, 2023).
@@
This paper delves into the creation of these student-centric guidelines, rendering them into clear instructions for LLM use. More specifically, we evaluate how various guidance strategies—including direct instruction, example-based teaching, metacognitive questioning, and worked examples—affect students’ performance, self-confidence, trust in the chatbot, and their perceptions of its efficacy.

### v3

@@ Abstract.
large language models, tutoring systems, Human-AI collaboration, artificial intelligence in education, collaborative learning with AI, transparency.
copyright: acmlicensedjournal: PACMHCIjournalyear: 2024journalvolume: 8journalnumber: CSCW2article: 499publicationmonth: 11doi: 10.1145/3687038ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Empirical studies in collaborative and social computingccs: Applied computing Computer-assisted instructionccs: Social and professional topics Computing education
@@ 1. Introduction
To address this issue, previous research on Intelligent Tutoring Systems (ITS) and Intelligent Teaching Assistants (ITA) has pioneered adaptive learning experiences and personalized instructional feedback (Weitekamp et al., 2020; Anderson et al., 1995), increasing the interest of CSCW community to the role that conversational agents can play in supporting the education process (Richards Maldonado et al., 2023; Wu et al., 2021; Breideband et al., 2023; Wardrip et al., 2013; Liu et al., 2020; Sobel et al., 2017; Davis et al., 2018). The rapid emergence of Large Language Models (LLMs) in recent years has created even more promising opportunities for technological support of personalized attention needs of students (Zhou et al., 2022; Sejnowski, 2023). These models, with their ability to comprehend and generate human-like text, have the potential to reshape the landscape of personalized education. Consequently, researchers started exploring the potential advantages of the development of LLM-based agent tutors as potential surrogates for human teaching assistants (Rahman and Watanobe, 2023; Dergaa et al., 2023; Baidoo-Anu and Owusu Ansah, 2023; Kumar et al., 2023b). Indeed, the dialogic nature of LLM-based agents, their ability to adapt to individualized learning patterns, and their ease of integration into diverse educational contexts make them a potential boon for educational frameworks (Farrokhnia et al., 2023). Although learning environments vary between subjects and educational levels, the versatility of LLMs allows their integration into these specialized contexts, providing a unified platform for diverse educational needs.
@@
However, although there is a considerable volume of existing research on LLM applications, a noticeable gap remains when applied to education. For example, a large body of pedagogical research (Keil and Wilson, 2000; Borko et al., 1997; Shepard, 2001; Van Driel et al., 2001; Chi et al., 2001) suggests that the success of education and tutoring is significantly defined by the communication strategies employed by an instructor. Thus, the pedagogical and interaction effects are inherently interconnected, and discerning their nuanced impacts is essential for the effective design and implementation of LLM tutors. However, while LLM-based systems can support different interaction strategies (Lee et al., 2022; Kasneci et al., 2023), it remains unclear which of these strategies are most beneficial for an agent to assume when interacting with students. For example, how could these strategies influence the students’ reception of the assistance, self-assessment, and their subsequent performance? The matter is further complicated by evidence that the dynamics of social interactions might differ for human-human and human-agent communications (Doyle et al., 2019; Clark et al., 2019), making the direct integration of the pedagogical knowledge on learning patterns into agent behavior even more challenging. Correspondingly, while there is a substantial body of literature on the impact of human-human teaching interaction strategies on learning outcomes, our understanding of how these dynamics translate to an LLM-learner context remains limited. Thus, the question is not only about the feasibility of these LLM-based tutors, but also about the ambiguity surrounding the effects of different interaction strategies on students. To be able to effectively adapt LLM-based agents for individual tutoring, we first need to understand what interaction strategies should be implemented in these systems for optimal learning support.
@@
What influence do various guidance strategies have on enhancing or reducing learners’ confidence in their ability to use LLMs for problem solving at each stage of interaction?
@@
In what specific ways do different guidance methods shape learners’ self-confidence in problem solving across the various stages of interaction?
@@
To gain a deeper understanding of student interactions with LLMs in educational contexts, in this paper, we distinguish three key elements of student-agent interactions: the Guidance Received, which refers to the directions or suggestions students obtain for using LLMs; the Learner Approach, indicating whether learners first consult the LLM or attempt problem solving independently; and the LLMs’ Response, emphasizing the importance of dialogue quality with LLM. These components are intertwined, with shifts in one potentially influencing the others. We then empirically investigated these influences through two user studies. In both studies, we explored four variations of the first element, using four distinct guidance strategies: list-based suggestions, example-based instruction, metacognitive questioning (making learners think about their use of LLMs before actual use) and a problem-solving-then-refinement approach (making learners solve a problem first and then refine it with LLMs). These four guidance strategies manifested in two distinct student approaches (second element): directly consulting the LLM or relying on self-first problem solving. Finally, we manipulated the third element – LLMs’ Response – through pre-interaction manipulations (Brown et al., 2020; Kumar et al., 2022).
@@
Although there is previous research on student-AI interaction, LLMs seem to provide different levels of interaction experience and convincingness, even in the presence of errors. Thus, studying the multidimensional picture of student perceptions and their dynamics and analyzing them in a formative experimental setup, combining preliminary quantitative measures as well as extensive qualitative feedback is needed at this initial stage. Our formative study in an undergraduate computer science classroom ($N=145$) tasked students with solving an assignment while receiving help from an LLM-based chatbot. We explored the impact of our four guidance strategies, and the resulting two approaches, on the student’s performance (in the assignment, as well as in the final exam 2 weeks after the assignment), system perception (helpfulness, confidence in LLM’s response, tolerance of mistakes, willingness to interact again), and their self-assessment (change in their self-efficacy, confidence in their answers). We found a reduction in unrelated student queries to LLM with structured guidance such as List of Suggestions, Metacognitive Questioning, and Solve then refine with LLM approaches, while Example based instruction paradoxically increased such queries. Furthermore, certain types of guidance promoted deeper engagement with the material, evidenced by fewer instances of verbatim question copying and increased rephrasing, highlighting the impact of guidance on the quality and focus of student-tutor interactions in LLM environments. As we discuss further, from a statistical point of view, this setting is traditional in its inherent limitations of sample size by course enrollment, student dropout, and selective engagement. It allows us to use it only to gather formative data for different conditions. In addition, exposing students to experimental stimuli allows us to gather qualitative feedback specific to different stimuli.
@@
Informed by findings from formative deployment, we conducted a controlled study on Prolific (Palan and Schitter, 2018) in which crowdworkers (N=356) were asked to solve math problems at the elementary and high school level with the help of an LLM chatbot. This math knowledge level was selected to ensure a standardized baseline of complexity and familiarity for a broad adult audience, allowing for a more controlled and consistent assessment of the LLM chatbot’s tutoring capabilities. Besides manipulating the different forms of guidance given to students, the online controlled setting allowed us to compare a pre-prompted LLM with an unprompted LLM. Through a 2 (LLM Chatbot: Prompted vs. Unprompted) x 5 (Guidance strategy: G1-List of Suggestions vs. G2-Example-based Instruction vs. G3-Metacognitive Questioning vs. G4-Solve then refine with LLM vs. No Instruction), we explored the impact of interaction strategies on the learners’ performance and their confidence (in LLM’s responses, ability to use LLMs, and their Math-solving ability) across the different stages of interaction (before receiving guidance, immediately after receiving guidance, and after task completion). We did not find statistically significant differences in Learner + AI task performance across different Guidance Types and LLM Types (prompted vs. unprompted), and we observe interesting patterns in confidence in the ability to use LLMs to solve math problems and trust in LLM responses over different stages of interaction. The guidance briefly boosts confidence in the ability to use LLM, but both measures significantly drop after learners attempt to solve real tasks using LLM. This might suggest a partially negative interaction experience in solving tasks, as well as the need for additional learning support for LLM use in the particular context.
@@ 2.1. Understanding Learner-LLM Interactions in Real-World Settings
With LLM-based conversational systems like ChatGPT (cha, [n. d.]), Bard (bar, [n. d.]), and BingChat (bin, [n. d.]) already out in the wild and used daily by hundreds of thousands of learners, educators and researchers face an urgent need to understand how learners interact with these systems in real-world educational settings. They are both thrilled (Rahman and Watanobe, 2023; Dergaa et al., 2023; Baidoo-Anu and Owusu Ansah, 2023) and concerned (Rahman and Watanobe, 2023; Dergaa et al., 2023) about the integration of LLM-based chatbots into diverse learning environments as we navigate the uncertainties regarding the impact of this rapidly evolving technology on pedagogy.
The growing body of recent CSCW and HCI literature on the incorporation of LLM-based conversational agents in education has explored various ways to enhance chatbots to assist learners in different ways, such as using them as persuasive agents to target behavior change through self-reflection (Li et al., 2023), making them more goal-aware (Deng et al., 2023), proactive (Liao et al., 2023), and credible (Candello et al., 2023), and customizing them to take on different roles to facilitate different discussion patterns and system thinking (Nguyen, 2023). However, while such works can provide valuable preliminary insights into how we can design chatbots, they often do not investigate learner interactions in real-world settings. Given that learners are already interacting with widely available and powerful LLMs through tools like ChatGPT, it has become increasingly important to understand how learners are interacting with these tools. In this work, we offer empirical insights of classroom and crowdworker interactions with LLMs from two field studies that contribute to this understanding, and offer design recommendations that can inform the design of future LLM-based chatbot implementations.
@@ 2.2. Leveraging Guidance Strategies to Steer Interactions Toward Learning
In conjunction with investigating learner-LLM interactions, we also need to understand how to guide students in ways that support their learning rather than impede it. Guidance plays a crucial role in helping users navigate how to best interact with chatbots, but the efficacy of specific forms of guidance remains underexplored (Wu et al., 2021).
@@
The CSCW community has begun to explore different aspects of this important space. For example, Wu et al. (Wu et al., 2021) studied user preferences for two different types of guidance (Example-Based and Rule-Based) at four different timings of guidance (Service-Onboarding, Task-Intro, After Failure and Upon Request) based on data from 24 participants, and found that users preferred Example-Based guidance, and wanted to see the guidance as part of Task-Intro. Zhu et al. (Zhu et al., 2022) built a chatbot prototype that provided task-based instructions to novice workers, and offered results from a pilot study with 7 participants that indicated that back-and-forth conversations could help novice workers follow instructions, and perceive those instructions as more actionable. Our work contributes to this body of literature on how different guidance strategies impact learner perceptions and interactions with LLM-based chatbots. A distinguishing feature of our work from existing preliminary studies is that we deployed our experiments in two distinct learner groups with comparatively larger sample sizes (145 students in an undergraduate CS classroom and 356 crowdworkers from Prolific). We study the specific effects of four guidance strategies (providing instructions, presenting examples, using metacognitive questioning, and encouraging initial problem solving with subsequent refinement).
@@
To inform our selection of guidance strategies, in addition to looking at the types of guidance that have been explored in preliminary work within the CSCW community, we also consulted previous pedagogical literature on various forms of instructional guidance (Bonawitz et al., 2011), and chose to study direct instruction (White, 1988), example-based instruction (Wittwer and Renkl, 2010), and metacognitive strategies (Livingston, 2003; Şen, 2009)—such as posing reflective questions (Ellis et al., 2014; Alt and Raichel, 2020), due to their documented effectiveness in prior research (Gersten et al., 1986; Adams and Engelmann, 1996; Zendler and Klein, 2018; Anderson et al., 2007; Van Gog and Rummel, 2010; Palincsar, 1986; Bourner, 2003).
@@
To understand why guidance is of particular relevance to LLM-based chatbot design, we turn to recent research on chatbot interactions of novice AI users that show prompt-writing can be deceptively difficult. The response from LLM-based chatbots is primarily steered by text-based prompts (Wu et al., 2022; Zhou et al., 2022). However, recent research has indicated that crafting effective prompts is challenging (Zamfirescu-Pereira et al., 2023), particularly for those without a deep understanding of AI (Abdellatif et al., 2020; Zamfirescu-Pereira et al., 2023; Fiannaca et al., 2023). The conversational interface of these chatbots mimics human interaction (Smestad and Volden, 2019), potentially misleading students into thinking that prompt-writing is as straightforward as talking to humans (Hofstadter, 1995). This gap between perceived simplicity and the actual complexity of effective prompt-writing can breed overconfidence. However, when the chatbot fails to respond as expected, students’ trust in the AI system (Denny et al., 2023a) can wane, leading to frustration (Pérez et al., 2020). In response to these challenges, informal communities like the r/aipromptprogramming and r/ChatGPT subreddits have emerged to share best practices for prompt-writing. Simultaneously, researchers are starting to establish formal guidelines (Denny et al., 2023b; Liu and Chilton, 2022; Ekin, 2023), but given the nascent nature of this field, much remains to explore. To improve student success with LLM, it is necessary to convert these guidelines into a more digestible format, thereby improving LLM literacy (Zamfirescu-Pereira et al., 2023; Yang, 2023).
@@
This paper investigates the creation of these student-centric guidelines, translating them into clear instructions for LLM use. More specifically, we evaluate how various guidance strategies—including direct instruction, example-based teaching, metacognitive questioning, and worked examples—affect students’ performance, self-confidence, trust in the chatbot, and their perceptions of its efficacy.